determine the general antiderivatives of the following functions f(x)=(x+1)100

1
Expert's answer
2022-05-22T23:48:06-0400

"\int f(x)dx=\int (x+1)^{100}dx=\frac {(x+1)^{101}}{101}+C, C\isin R"